OpenStudy (anonymous):

Is the ordered pair (7,14) a solution of the inequality 3y-4x>7

OpenStudy (unklerhaukus):

if x= 7, y=14, 3y-4x becomes 3x14-4x7 =42-28 =14 which is greater than 7

OpenStudy (anonymous):

so I am right is saying my answer is yes

OpenStudy (unklerhaukus):

The ordered pair is a solution to the equality. (YES)

OpenStudy (unklerhaukus):

*the inequality